Formula for a compound with A at corners and B at alternate edges....??

Grade:12

1 Answers

Vikas TU
7 years ago
R u talking about the solid states chapter?
and what is solid state but?
cubical.?
ok assuming cubical, we get the answer as:
 
A = 1/8
B= ¼ * (12 -3) = 9/4
Formulae : A1/8B9/4
Or
8(AB18)